# Energy consumption of code small language models serving with runtime engines and execution providers

### v1

@@ 5 Results
CPU usage: The visual data and median values indicate that configurations using the CPU execution provider generally exhibit higher CPU usage than those employing the CUDA execution provider, which aligns with expectations given the operational differences between the two. For CUDA execution providers, CPU bottlenecks are minimal since the majority of computations are offloaded to the GPU. However, when using CPU execution provider, as it is expected that all the operations are executed by the CPU, selecting the serving configuration may impact CPU usage.
The $<$TORCH, CUDA$>$ configuration consistently demonstrated the lowest CPU usage among all configurations, highlighting its efficiency in resource utilization. This configuration achieves an effective balance between GPU and CPU usage, optimizing the model operations for improved performance.
@@
The $<$TORCH, CUDA$>$ configuration consistently demonstrated the lowest RAM usage, ranging from 4.77% to 11.45% in 9 out of 12 models. This highlights its efficiency in resource utilization, as it achieves a balance between computational performance and minimal RAM consumption. TORCH configurations, in general, exhibit lower RAM usage compared to other runtime engines.
In contrast, $<$ONNX, CPU$>$ and $<$OV, CPU$>$ configurations exhibited significantly higher RAM usage, with $<$ONNX, CPU$>$ ranging from 9.96% to 30.16% and $<$OV, CPU$>$ from 8.85% to 20.88% in 10 out of 12 models. These findings indicate that optimized runtime engines like ONNX and OV demand more system resources due to their memory management strategies. While they improve energy efficiency and execution time in some scenarios, this often comes at the cost of increased RAM usage. Furthermore, in CUDA execution provider configurations, ONNX exhibits higher RAM usage but performs worse in terms of both time and energy efficiency compared to TORCH, further underscoring TORCH’s superior balance of performance and resource utilization.
GPU usage: results show that the configuration $<$TORCH, CUDA$>$ reaches the highest levels of GPU utilization (from 21.08% up to 86.98% ), leveraging available GPU resources effectively. Conversely, the $<$ONNX, CUDA$>$ configuration shows the lowest GPU usage (from 18.07% up to 75.49% ) among the CUDA execution provider options, suggesting it may not exploit the computational potential of the GPU as efficiently as $<$TORCH, CUDA$>$. This trend is also reflected in the aggregated results across the three groups.
GPU memory utilization: Our GPU utilization memory results reveal that $<$TORCH, CUDA$>$ configurations exhibited the highest memory utilization (from 11.57% up to 83.59%, all models), effectively leveraging GPU capabilities for intensive memory operations. In contrast, $<$ONNX, CUDA$>$ configurations showed the lowest memory utilization (from 2.34% up to 27.33%), suggesting either conservative use or inefficiencies in exploiting GPU VRAM operations compared to configurations using $<$TORCH, CUDA$>$.
@@
GPU used memory:Regarding GPU allocated memory (GPU used memory), the results show varying patterns among CUDA execution provider configurations. For 7 out of 12 models, the $<$TORCH, CUDA$>$ configuration exhibited higher levels of memory allocation, while the $<$ONNX, CUDA$>$ configuration showed higher memory allocation for the remaining 5 models. Specifically, $<$TORCH, CUDA$>$ configurations ranged from 12.01% to 59.16%, whereas $<$ONNX, CUDA$>$ configurations ranged from 5.12% to 71.87%. These findings highlight differences in memory allocation strategies and the relative efficiency of Torch in allocating GPU memory.
Answer to RQ3: Computing-resource utilization is significantly influenced by the serving configuration, with OV or ONNX within CPU configurations improving energy or time efficiency but increasing RAM demands. In contrast, CUDA configurations, particularly $<$TORCH, CUDA$>$, show more efficient resource utilization, achieving higher GPU utilization and better memory allocation, reducing the load on CPU and RAM.
@@ 6 Discussion and implications
Energy consumption. In our analysis, we observe that configurations utilizing the CUDA execution provider consistently consume less energy compared to those utilizing the CPU execution provider. This significant difference underscores the inherent efficiencies of CUDA execution provider, which enables hardware-accelerated computation on NVIDIA GPUs, effectively managing parallel processing tasks in code generation models.
Interestingly, ONNX runtime configurations show greater optimization when paired with CPU execution provider compared to CUDA execution provider, when compared to other runtime engines under the same execution provider. This enhanced performance with CPU execution provider could stem from the specific libraries used (such as Transformers and Optimum HuggingFace’s libraries [59]) or the particular demands of code generation SLMs. It can be resource-intensive for ONNX configurations because of the overhead involved in mapping different subgraphs to suitable execution providers and the synchronization of inputs [63]. However, this complex binding process could yield benefits in scenarios involving varied batch sizes or other types of deep learning tasks, where such dynamic resource allocation is advantageous. Further exploration of this type of runtime engine across various deep learning tasks could provide valuable insights for software engineers.
The findings from our study highlight crucial implications for deploying deep learning models, particularly in how different configurations affect energy consumption and performance. The superior energy efficiency of CUDA execution provider configurations suggests significant potential for cost savings and reduced environmental impact, making GPU-accelerated environments preferable for intensive parallel processing tasks. Furthermore, the trade-offs observed with ONNX’s varied optimization across execution providers, underline the importance of aligning runtime engines with the appropriate hardware to balance energy efficiency with computational accuracy.
Execution time. CUDA execution provider configurations consistently outpace those using CPU execution provider, demonstrating the substantial advantages of leveraging NVIDIA GPUs for parallel processing tasks. This performance disparity is consistent with findings from other studies, further confirming the superior processing speeds of GPUs.
@@
The ONNX runtime engines shows a notable difference in optimization between CPU execution provider and CUDA execution provider configurations. It performs well under CPU execution provider, likely due to tailored optimizations for CPU usage, but does not translate these benefits as effectively to CUDA execution provider, when comparing to the other runtime engines. We find this counterintuitive as GPU implementations are usually faster than CPU, it could be to [64] or that CUDA execution provider in ONNX runtime does not fully support CUDA kernels used for the SLMs [65]. Due to the speed of development of deep learning research and development, we expect this type of optimized runtime engines are in general a better option for production-like deep learning systems.
@@
In their study, Bekhelifi et al. [63] found that configurations using $<$ONNX, CPU$>$ outperformed those using $<$OV, CPU$>$ in terms of speed for Brain-Computer Interface applications. However, our results contrast this finding, as $<$OV, CPU$>$ proved faster in our experiments. This discrepancy highlights how the performance of optimized runtime engines can vary significantly across different deep learning tasks. We recommend further investigation into how these engines perform across a broader range of deep learning systems to better understand their task-specific efficiencies and limitations.
@@
Computing-resource utilization. Configurations using CPU execution provider typically encounter bottlenecks in CPU and RAM usage. This underscores the need for careful selection and optimization of runtime engines to avoid overloading the CPU, as computational tasks in this execution provider heavily rely on RAM. In contrast, CUDA execution provider configurations benefit from the additional use of VRAM, distributing the memory load more effectively between system RAM and GPU memory. For instance, ONNX and OV configurations, while energy-efficient in CPUEP, require substantial memory (RAM and/or GPU memory) to handle inference, Using lower precision can help manage GPU memory when using CUDA execution provider [65].
@@
Within the CUDA execution provider configurations, significant statistical differences were observed in GPU utilization and memory efficiency. The $<$TORCH, CUDA$>$ configuration consistently demonstrated the highest GPU usage, GPU memory utilization, and GPU used memory, highlighting its ability to effectively harness GPU resources for computationally intensive tasks. In contrast, the $<$ONNX, CUDA$>$ configuration showed comparatively lower GPU utilization, suggesting potential inefficiencies in its optimization for the CUDA execution provider or its compatibility with specific code generation SLMs. Despite its higher GPU resource utilization, $<$TORCH, CUDA$>$ proved to be more efficient overall, as it consumed less CPU and RAM compared to $<$ONNX, CUDA$>$, indicating a balanced and resource-efficient approach to managing deep learning inference workloads.
@@
Torchscript runtime. As outlined in Section 5, the TorchScript runtime engine was also included in our experiments. However, we excluded its results from the analysis due to the error-prone outputs it produced. Instead, we investigated the causes of these issues, focusing on underlying factors rather than setup problems.
@@
As a runtime engine, TorchScript’s primary mode of operation in this study was through Trace mode, which inherently supports a broader range of Python features by operating beneath the Python layer. This mode records all PyTorch operations executed with a representative input to create an optimized TorchScript Intermediate Representation (IR). While this mode is advantageous for accommodating all Python features and ensuring consistent execution paths, it is notably error-prone as it does not capture dynamic control flows inherent in Python, it only supports functions with a single path, potentially leading to inaccuracies in conditions where execution paths vary [66, 67, 68].
@@
TorchScript’s ability to fuse operations into larger computational blocks offers potential advantages by reducing memory reads and writes, thereby optimizing certain computational tasks. Additionally, intermediate results can be cached directly in registers or shared memory, minimizing the need for repeated memory access [69, 70]. However, despite these memory optimizations, overall memory usage remains high, indicating that the benefits do not necessarily translate into lower memory consumption. This limitation, combined with TorchScript’s inability to effectively manage dynamic execution paths, diminishes its potential energy and speed advantages. These findings highlight the critical need to align runtime engine capabilities with the specific demands and infrastructure of the deployment environment to achieve optimal performance and computing-resource utilization.
@@
If a static graph exported shows good responses, it could be the best option for a production environment, however it must be really well tested and it is time-exhaustive. There are some options being built that can potentially eliminate this drawbacks [66, 67].
@@
    Energy efficiency optimization: CUDA execution provider configurations, particularly $<$TORCH, CUDA$>$, are the most energy-efficient for code generation tasks. CPU execution provider configurations, when paired with ONNX runtime engines, offer a balanced alternative, optimizing resource usage without the need for GPU hardware, making them ideal for energy-aware, non-GPU environments.
@@
    Execution time optimization: CUDA execution provider configurations consistently deliver faster performance across the board. TORCH under CUDA execution provider offers the fastest execution times. In CPU configurations, OV provides the best execution time enhancements.
@@
    Computing-resource utilization: In CPU execution provider configurations, if RAM is a constraint, $<$TORCH, CPU$>$ offers a more balanced option, though it may not be optimal for energy efficiency or execution time. For better overall resource management, $<$TORCH, CUDA$>$ within CUDA execution provider configurations stands out, showing better resource management in overall(i.e. leveraging GPU resources while reducing workload to CPU, and RAM memory).
Optimizing LMs inference. Scaling up the size of language models has consistently been shown to improve their accuracy [71]. However, this scaling comes with substantial computational costs, memory demands, and environmental concerns, posing significant challenges for the practical deployment of LMs. This study, along with prior research [16], highlights the critical role of serving configurations in improving energy efficiency and performance, demonstrating how their selection can significantly impact these metrics.
The increasing adoption of LMs in real-world applications highlights the critical role of system-level decisions. In addition to runtime engines and execution providers, serving infrastructures such as DL-specific software (e.g., TorchServe) and end-to-end cloud services (e.g., AWS SageMaker) represent system-level optimizations that can enhance energy and performance efficiency without compromising model accuracy [16, 72]. These decisions align with the software engineering perspective, where employing green software tactics can effectively optimize deep learning systems throughout their lifecycle [73].
@@
7 Limitations & threats to validity

### v2

@@ 5 Results
We profiled computing-resource utilization using described profiling tools, capturing explained metrics. VRAM (GPU memory) was only utilized in configurations involving the CUDA execution provider. In CPU-only configurations, VRAM usage remained consistently negligible or zero, as expected. These results confirm the distinction in memory handling strategies across execution providers.
CPU usage: The visual data and median values indicate that configurations using the CPU execution provider generally exhibit higher CPU usage than those employing the CUDA execution provider, which aligns with expectations given the operational differences between the two. For CUDA execution providers, CPU bottlenecks are minimal since the majority of computations are offloaded to the GPU. Conversely, CPU-based configurations concentrate the workload entirely on the processor and system memory, resulting in higher RAM and CPU usage.
The $\langle TORCH,CUDA\rangle$ configuration consistently showed the lowest CPU usage among all configurations, indicating that it effectively shifts computation to the GPU wherever possible. By reducing reliance on the CPU and leveraging GPU capabilities for parallel operations, this setup optimizes computing-resource utilization.
@@
The $\langle TORCH,CUDA\rangle$ configuration consistently demonstrated the lowest RAM usage, ranging from 4.77% to 11.45% in 9 out of 12 models. These percentages correspond to the cases where $\langle TORCH,CUDA\rangle$ ranked lowest for RAM usage among all GPU-based configurations. In the remaining 3 models, a different GPU-based configuration achieved the lowest RAM usage. This highlights its efficiency in computing-resource utilization, as it achieves minimal energy and execution time, and minimal RAM consumption. TORCH configurations, in general, exhibit lower RAM usage compared to other runtime engines.
In contrast, $\langle ONNX,CPU\rangle$ and $\langle OV,CPU\rangle$ configurations exhibited significantly higher RAM usage, with $\langle ONNX,CPU\rangle$ ranging from 9.96% to 30.16% and $\langle OV,CPU\rangle$ from 8.85% to 20.88% in 10 out of 12 models. These findings indicate that optimized runtime engines like ONNX and OV demand more system resources due to their memory management strategies. While they improve energy efficiency and execution time in some scenarios, these benefits often involve higher RAM usage. In scenarios with limited computational resources, increased memory requirements can indeed become a significant concern, outweighing gains in execution time or energy savings. Furthermore, in CUDA execution provider configurations, ONNX exhibits higher RAM usage but performs worse in terms of both time and energy efficiency compared to TORCH, further underscoring TORCH’s superior results of performance in all three dependent metrics.
GPU usage: results show that the configuration $\langle TORCH,CUDA\rangle$ reaches the highest levels of GPU utilization (from 21.08% up to 86.98%), indicating that a substantial portion of the available GPU compute capacity is actively used during inference. Conversely, the $\langle ONNX,CUDA\rangle$ configuration shows the lowest GPU usage (from 18.07% up to 75.49% ) among the CUDA execution provider options, suggesting it may not exploit the computational potential of the GPU as efficiently as $\langle TORCH,CUDA\rangle$. This trend is also reflected in the aggregated results across the three categories.
GPU memory utilization: Our GPU utilization memory results reveal that $\langle TORCH,CUDA\rangle$ configurations exhibited the highest memory utilization (from 11.57% up to 83.59%, all models), effectively leveraging GPU capabilities for intensive memory operations. In contrast, $\langle ONNX,CUDA\rangle$ configurations showed the lowest memory utilization (from 2.34% up to 27.33%), suggesting conservative number of GPU VRAM operations compared to configurations using $\langle TORCH,CUDA\rangle$.
@@
GPU used memory: Regarding GPU allocated memory (GPU used memory), the results show varying patterns among CUDA execution provider configurations. For 7 out of 12 models, the $\langle TORCH,CUDA\rangle$ configuration exhibited higher levels of memory allocation, while the $\langle ONNX,CUDA\rangle$ configuration showed higher memory allocation for the remaining 5 models. Specifically, $\langle TORCH,CUDA\rangle$ configurations ranged from 12.01% to 59.16%, whereas $\langle ONNX,CUDA\rangle$ configurations ranged from 5.12% to 71.87%. These findings highlight differences in memory allocation strategies.
Answer to RQ3: Computing-resource utilization is significantly influenced by the serving configuration. CPU-based configurations (particularly with OV or ONNX) exhibit higher RAM usage, reflecting increased memory demands. In contrast, CUDA configurations, particularly $\langle TORCH,CUDA\rangle$, show more efficient computing-resource utilization, achieving higher GPU utilization and better memory allocation, reducing the load on CPU and RAM.
@@
((a)) Energy consumption [J]
((b)) Energy consumption [J]
((c)) Execution time [s]
((d)) Execution time [s]
Figure 6: Example of energy consumption and execution time for model pythia-410m before model size category aggregation: CUDA execution provider (left) and CPU execution provider (right) measured by Energibridge and Nvidia-smi
((a)) CPU usage [%]
((b)) CPU usage [%]
((c)) RAM usage [%]
((d)) RAM usage [%]
Figure 7: Example of computing-resource utilization for model pythia-410m before model size category aggregation: CUDA execution provider (left) and CPU execution provider (right) measured by Energibridge
((a)) GPU utilization [%]
((b)) Utilization GPU memory utilization [%]
((c)) Used GPU memory utilization [%]
Figure 8: Example of computing-resource utilization for model pythia-410m before model size category aggregation: CUDA execution provider configurations measured by Nvidia-smi
@@ 6 Discussion and implications
Energy consumption. In our analysis, we observe that configurations utilizing the CUDA execution provider consistently consume less energy compared to those utilizing the CPU execution provider. This significant difference underscores the inherent efficiencies of CUDA execution provider, which enables hardware-accelerated computation on NVIDIA GPUs, effectively managing parallel processing tasks in code generation SLMs.
Interestingly, ONNX runtime configurations show greater optimization when paired with CPU execution provider compared to CUDA execution provider, when compared to other runtime engines under the same execution provider. This enhanced performance with CPU execution provider could stem from the specific libraries used (such as Transformers and Optimum HuggingFace’s libraries [74]) or the particular demands of code generation SLMs. It can be resource-intensive for ONNX configurations because of the overhead involved in mapping different subgraphs to suitable execution providers and the synchronization of inputs [78]. However, this complex binding process could yield benefits in scenarios involving varied batch sizes or other types of deep learning tasks, where such dynamic resource allocation is advantageous.
Although $\langle TORCH,CUDA\rangle$ registers the lowest total energy in our experiments, two runtime engine artifacts make its lead over $\langle ONNX,CUDA\rangle$ look larger than it might truly be. First, limited kernel coverage: several decoder-specific operators lack CUDA implementations in the evaluated ONNX Runtime release, so they fall back to the CPU execution provider, raising host load and data transfers [79, 80]. Second, sub-optimal I/O binding (IOBinding): every operator executed on the CPU incurs host-device and device-host copies, further incrementing energy use [81]. The runtime engine logs (“24 Memcpy nodes added”, “Some nodes not assigned to the preferred EP”) confirm both the extra transfers and CPU fall-back for some nodes. Practitioners should therefore enable IOBinding, review verbose EP-placement logs and verify that critical nodes execute on the intended hardware before drawing conclusions. Further exploration of this type of runtime engine across various deep learning tasks could provide valuable insights for software engineers.
The findings from our study highlight crucial implications for deploying deep learning models, particularly in how different configurations affect resource utilization. The superior energy efficiency of CUDA execution provider configurations suggests significant cost savings and reduced environmental impact, making GPU-accelerated environments preferable for intensive parallel processing tasks. Furthermore, the trade-offs observed with ONNX’s varied optimization across execution providers, underline the importance of aligning runtime engines with the appropriate hardware to reduce energy consumption, without requiring changes to the model architecture or applying model-level optimizations such as model compression [82].
Execution time. CUDA execution provider configurations consistently outpace those using CPU execution provider, demonstrating the substantial advantages of leveraging NVIDIA GPUs for parallel processing tasks. This performance disparity is consistent with findings from other studies, further confirming the superior processing speeds of GPUs [30].
@@
The ONNX runtime engine shows a significant difference in optimization between CPU execution provider and CUDA execution provider configurations. It performs well under CPU execution provider, likely due to tailored optimizations for CPU usage, but does not translate these benefits as effectively to CUDA execution provider, when comparing to the other runtime engines. We find this counterintuitive, given the common expectation that GPU implementations typically outperform CPU ones. We believe this may be due to two main factors: (i) the CUDA execution provider in ONNX runtime does not fully support CUDA kernels used for the SLMs [83], and (ii) our selected models or batch size may not have been complex enough to fully exploit GPU parallelism, or we may not have identified the most optimal CUDA-specific configuration within the used library [84]. These findings underscore the importance of aligning software optimizations with the characteristics of the target hardware. Due to the speed of development of deep learning research and development, we expect that this type of optimized runtime engines are, in general, a better option for real-world deep learning systems that require efficient resource utilization.
The median time advantage of $\langle TORCH,CUDA\rangle$ over $\langle ONNX,CUDA\rangle$ is also related to kernel available operators for CUDA execution provider and data transfers. Profiling showed that the remaining difference was driven by host–device memcpy nodes inserted by the ONNX optimizer and operations falling back on the CPU execution provider. In other words, the gap can be largely an artifact of kernel-support and data transfer overheads rather than an intrinsic limitation of ONNX Runtime. Our results favor $\langle TORCH,CUDA\rangle$ partly because experiments ran on a single NVIDIA GeForce RTX 4090, a GPU for which TORCH has targeted optimization due to its common use as deep learning framework. ONNX Runtime, by contrast, still lacked CUDA kernels for some decoder-centric kernels in the used library version. On other execution providers or future ONNX releases with full kernel parity and IOBinding well enabled, the ranking could invert [80]. We therefore recommend reviewing logs and verify optimizations like IOBinding are being used to exploit the faster kernels for the used hardware.
@@
In their study, Bekhelifi et al. [78] found that configurations using $\langle ONNX,CPU\rangle$ outperformed those using $\langle OV,CPU\rangle$ in terms of execution time for Brain-Computer Interface applications. For example, they observed that $\langle ONNX,CPU\rangle$ resulted in fewer backend calls compared to $\langle OV,CPU\rangle$ when profiling inference execution, suggesting that ONNX may execute a more streamlined inference. However, our results contrast with this finding, as $\langle OV,CPU\rangle$ proved to be faster in our experiments. This discrepancy highlights how the performance of optimized runtime engines can vary significantly across different deep learning tasks. Similarly, Pochelu et al. reported that OV consistently outperformed other CPU-based serving configurations for CNN workloads, while emphasizing that this choice should align with both application requirements and hardware characteristics [85].
We recommend further investigation into how these engines perform across a broader range of deep learning systems to better understand their task-specific efficiencies and limitations. Additionally, we recommend that software engineers assess inference performance within their specific deployment contexts and take advantage of the rapid experimentation capabilities offered by the tools in our replication package to identify the most suitable configuration for their needs.
Optimizing execution time frequently aligns with reduced energy consumption due to shortened computational workloads. However, this alignment is explicitly not universal. Certain configurations, notably $\langle OV,CPU\rangle$, optimized execution time yet showed increased energy demands compared to slower configurations like $\langle ONNX,CPU\rangle$, clearly demonstrating that faster execution can occasionally entail greater energy costs. Consequently, we explicitly advocate against using execution time as a sole proxy for energy efficiency. Instead, direct monitoring of energy consumption is recommended to avoid misinterpreting or overlooking discrepancies among the key metrics for the specific deep learning system.
@@
Computing-resource utilization. Configurations using CPU execution provider typically encounter bottlenecks in CPU and RAM usage. This underscores the need for careful selection and optimization of runtime engines to avoid overloading the CPU, as computational tasks in this execution provider heavily rely on RAM. In contrast, CUDA execution provider configurations benefit from the additional use of VRAM, distributing the memory load more effectively between system RAM and GPU memory. For instance, ONNX and OV configurations, while energy-efficient in CPU-based configurations, require substantial memory (RAM and/or GPU memory) to handle inference. Using lower precision can help manage GPU memory when using CUDA execution provider [83].
@@
Within the CUDA execution provider configurations, significant statistical differences were observed in GPU utilization and memory efficiency. The $\langle TORCH,CUDA\rangle$ configuration consistently demonstrated the highest GPU usage, GPU memory utilization, and GPU used memory, indicating a greater reliance on GPU resources to perform inference computations. In this context, higher GPU utilization reflects the configuration’s ability to effectively offload workload from the CPU to the GPU, which is particularly desirable for computationally intensive tasks. In contrast, the $\langle ONNX,CUDA\rangle$ configuration showed comparatively lower GPU utilization, suggesting potential inefficiencies in how it distributes computation or interacts with specific code generation SLMs. Despite its higher GPU resource usage, $\langle TORCH,CUDA\rangle$ proved more efficient overall, as it consumed less CPU and RAM and achieved superior energy and execution time metrics. This suggests a balanced and resource-conscious allocation strategy that fully leverages available GPU capabilities.
@@
JIT runtime. As outlined in Section 5, the JIT runtime engine was also included in our experiments. However, we excluded its results from the analysis due to the error-prone outputs it produced. Instead, we investigated the causes of these issues, focusing on underlying factors rather than setup problems.
@@
As a runtime engine, JIT’s primary mode of operation in this study was through Trace mode, which inherently supports a broader range of Python features by operating beneath the Python layer. This mode records all PyTorch operations executed with a representative input to create an optimized TorchScript Intermediate Representation (IR). While this mode is advantageous for accommodating all Python features and ensuring consistent execution paths, it is notably error-prone as it does not capture dynamic control flows inherent in Python, it only supports functions with a single path, leading to inaccuracies in conditions where execution paths vary [86, 87, 88].
@@
JIT’s ability to fuse operations into larger computational blocks offers advantages by reducing memory reads and writes, thereby optimizing certain computational tasks. Additionally, intermediate results can be cached directly in registers or shared memory, minimizing the need for repeated memory access [89, 90]. However, despite these memory optimizations, overall memory usage remains high, indicating that the benefits do not necessarily translate into lower memory consumption. This limitation, combined with JIT’s inability to effectively manage dynamic execution paths, diminishes its potential energy and speed advantages. These findings highlight the critical need to align runtime engine capabilities with the specific demands and infrastructure of the deployment environment to achieve optimal performance and computing-resource utilization.
@@
If a static graph exported shows good responses, it could be the best option for a production environment, however it must be really well tested and it is time-exhaustive. There are some options being built that can eliminate this drawbacks [86, 87].
Despite its exclusion from our final set of serving configurations, JIT remains an interesting runtime engine thanks to its operation-fusion capabilities and reduced overhead in certain scenarios. Specifically, when a model’s inputs remain stable or do not require complex branching and thus do not rely on newly generated outputs being fed back into the input (as often happens in code generation) JIT can deliver notable performance gains [34].
Our follow-up investigation revealed that the Trace mode we used struggles with the dynamic control flows central to code generation tasks. If an input prompt triggers a path beyond the initially traced computational graph, the resulting outputs may be incomplete or incorrect. This limitation reflects JIT’s reliance on a fixed execution graph, which does not adapt well to unexpected input variations. These observations align with prior work indicating JIT’s lower robustness in settings where the execution path frequently changes [86, 87, 88].
While JIT remains an optimal solution in scenarios featuring minimal branching and stable inputs, it requires further refinement to reliably support more dynamic workloads, such as code generation.
@@
    Battery or environmentally constrained serving (e.g., edge inference): We recommend prioritizing energy consumption. CUDA execution provider configurations, particularly $\langle TORCH,CUDA\rangle$, are the most energy-efficient for code generation tasks. When GPU access is limited, CPU execution provider paired with ONNX runtime offers a balanced and low-energy alternative, making them ideal for energy-aware, non-GPU environments.
@@
    Latency constrained serving (e.g., real-time code completion in IDEs): We suggest prioritizing execution time. CUDA execution provider configurations consistently deliver faster performance across the board. Among these, $\langle TORCH,CUDA\rangle$ offers the fastest performance. In CPU configurations, OV provides the best execution time enhancements.
@@
    Computing-resource constrained serving (e.g., legacy systems): In CPU execution provider configurations, if RAM is a constraint, we suggest using $\langle TORCH,CPU\rangle$ as a more computing-resource balanced option, although it may not be optimal in terms of energy efficiency or execution time. For better overall resource management, $\langle TORCH,CUDA\rangle$ stands out among CUDA execution provider configurations, as it effectively leverages GPU resources while reducing the workload on the CPU and minimizing RAM usage.
This study contributes practical insights that can support software engineers in the design, deployment, and maintenance of resource efficient deep learning systems. By incorporating non-functional properties, such as energy consumption, execution time, and computing-resource utilization, into the evaluation of serving infrastructure, our methodology enables practitioners to make informed architectural decisions during system design and operation.
Specifically, engineers can replicate or adapt our experimental framework to analyze the resource utilization of their own deep learning inference under different infrastructure configurations. These insights can guide choices in runtime engines and execution providers, depending on the constraints and priorities of a given application context (e.g., energy efficiency, latency, or hardware limitations).
Moreover, our findings highlight the architectural impact of serving infrastructure components, which are often overlooked in early design phases. As shown in this study and in prior work [16], such components can significantly influence resource utilization related properties, underscoring the need to include them in architectural planning.
Finally, system-level monitoring of non-functional metrics can serve not only as a benchmarking tool but also as a continuous feedback mechanism in production environments. This supports ongoing system improvements and helps teams maintain alignment with operational and environmental goals over time.
Optimizing language models inference. While scaling up the size of language models has consistently been shown to improve their accuracy [91], it also introduces significant computational, memory, and environmental costs that can hinder practical deployment. In contrast, SLMs offer a more efficient alternative, particularly suited for code generation tasks, as they provide acceptable performance with lower resource demands. In this context, our study, along with prior research [16], emphasizes that optimizing serving configurations plays a central role in improving inference resource utilization. These system-level decisions can enhance resource utilization without modifying the model or relying on techniques such as model compression, offering a practical approach for the sustainable deployment of SLMs.
The increasing adoption of language models in real-world applications highlights the critical role of system-level decisions. In addition to runtime engines and execution providers, serving infrastructures such as DL-specific software (e.g., TorchServe) and end-to-end cloud services (e.g., AWS SageMaker) represent system-level optimizations that can enhance energy and performance efficiency without compromising model accuracy [16, 92]. These decisions align with the software engineering perspective, where employing green software tactics can effectively optimize deep learning systems throughout their lifecycle [93].
@@
7 Limitations and threats to validity